z(x+y)p+z(x-y)q=x^2+y^2 


1
Expert's answer
2021-06-14T19:17:49-0400

"\text{By lagrange method,}\\
\frac{dx}{z(x+y)}=\frac{dy}{z(x-y)}=\frac{dz}{x^2-y^2}\\
first,\\
\frac{dx}{z(x+y)}=\frac{dy}{z(x-y)}\\
\frac{dx}{x+y}=\frac{dy}{x-y}\\
xdx-ydy-xdy-ydx=0\\
xdx-ydy-d(xy)=0\\
\text{Integrating, we get}\\
x^2-y^2-2xy=c_1\\
Second,\\
\frac{xdx-ydy-zdz}{z(x+y)-z(x-y)-z(x^2+y^2)}=0\\
xdx-ydy-zdz=0\\
x^2-y^2-z^2=c_2\\
Therefore,\\
\text{solution is given by,}\\
F(x^2-y^2-2xy,x^2-y^2-z^2)=0"